OpenStudy (anonymous):

demoivre's theorem A.Write (-2-2i)^7 in a+bi form B. Find the five fith roots of the complex number 1/2 -(sqrt3/2i).

OpenStudy (anonymous):

first write \(-2-2i\) in the form \(r(\cos(\theta)+i\sin(\theta))\) \[r=\sqrt{a^2+b^2}=\sqrt{2^2+2^2}=\sqrt8\]

OpenStudy (anonymous):

to find \(\theta\) if you plot the number it will be more or less obvious what the angle is